# Deterministic Legal Retrieval: An Action API for Querying the SAT-Graph RAG

### v1

@@ 3.3.6 Introspection & Metadata Actions
    • Discussion: This action enables an agent to be aware of its own knowledge boundaries. It should be used to proactively validate a user’s temporal query before attempting retrieval, allowing for graceful handling of requests for dates outside the known history and improving the overall user experience.
  2. 22.
@@
    • Discussion: This action provides the canonical list of values that can be used in the action_types filter of other functions like summarizeImpact. It allows a client application to dynamically build filtering interfaces without hardcoding ontology-specific values.
  4. 24.
@@
    • Discussion: This action serves as the primary and most efficient entry point for an agent or user interface to explore the thematic ontology. By returning the full root objects, it allows a client to render the top-level categories without requiring subsequent API calls.
@@ 3.3.7 Architectural Rationale and Modeling Principles
Our framework is designed to handle various types of legislative state transitions with precision. A notable example is the handling of revocations. A recommended practice for maintaining a continuous version timeline is to model a revocation not just as the termination of the last valid Version, but also as the creation of a new, subsequent Version with a specific textual content, such as "(Revoked)". However, the API is flexible enough to model this differently. This flexibility is fully represented in the Action data model, which can cleanly represent an original creation (with only a produces_version_id), a simple repeal (with only a terminates_version_id), or a standard amendment (using both).
Discussion: Atomic Actions and Emergent Event Cohesion
A core design principle of the SAT-Graph is the atomicity of causal events. Our model defines an Action as the most granular unit of change possible to ensure maximum auditability. A key consequence is that a single, complex legislative event—such as the enactment of "Constitutional Amendment No. 99"—is not represented by a monolithic node. Instead, its total effect is captured by an emergent collection of many individual Action nodes, whose cohesion is established by their respective source_version_id properties all tracing back to the same source Work.
This design provides the best of both worlds. An agent can use a single, high-level API action to reconstruct the full scope of an event:
@@
This powerful aggregation enables high-level impact analysis ("What were all the effects of this amendment?"), while the atomic nature of each returned Action preserves the granularity needed for pinpoint-accurate causal tracing.
@@
This flexibility is directly exposed through the API, particularly via the aspects parameter in functions like searchTextUnits. All TextUnits, regardless of their source entity or aspect, are candidates for embedding and semantic retrieval. This creates a rich, multi-faceted search space where an agent can retrieve information based not just on the law’s formal content, but also on its context, metadata, and conceptual definitions in a single, unified query.
@@ 3.4 Practical Implementation Considerations
While the actions specified above define the conceptual and logical interface to the SAT-Graph, a production-grade implementation should include additional features for efficiency and robust operation.
@@
To mitigate network latency in scenarios requiring numerous lookups—such as reconstructing the state of a constitutional chapter with 50 articles at a specific date, which would otherwise require 50 sequential API calls—, such as resolving the point-in-time structure of a complex item or analyzing its full history, the API should provide batch-enabled versions of its core retrieval actions. Key batch actions include:
  1. 25.
@@
    • Discussion: This is a fundamental utility action used after any discovery process. When an agent obtains a list of ids from functions like ‘searchItems‘ or ‘enumerateItems‘, this batch action allows it to hydrate those references into full objects to access their labels, types, or hierarchical relationships without incurring the high cost of N+1 individual requests.
  2. 26.
    getBatchValidVersions(item_ids: list[ID], timestamp: date-time,
    policy: TemporalPolicy) $\rightarrow$ list[Version]
@@
      • policy: The temporal resolution policy to apply uniformly.
@@
    • Discussion: This action is essential for efficiently reconstructing the state of a complex hierarchical item at a specific point in time. It is designed to be used directly with the output of ‘enumerateItems‘, avoiding the "N+1 query problem" and drastically improving performance.
  3. 27.
    getBatchTexts(version_ids: list[ID], language: string,
    aspects?: list[string]) $\rightarrow$ list[TextUnit]
    • Description: A batch-optimized version of getTextForVersion. Given a list of Version identifiers, it efficiently retrieves their corresponding textual content.
@@
      • version_ids: A list of identifiers for the Versions of interest.
      • language: The language to retrieve.
      • aspects: (Optional) A list of textual aspects to retrieve for all Versions (e.g., [‘"canonical"‘, ‘"summary"‘]). Defaults to [‘"canonical"‘].
    • Returns: A list of the requested TextUnit objects.
    • Discussion: This action is critical for use cases that require inspecting the content of multiple versions, such as finding the exact point of a textual change within a long version history. It prevents the severe performance degradation that would result from iterating and calling getTextForVersion individually for each version.
  4. 28.
@@
    • Discussion: This action is essential for the "hydration" step of analytical queries. After an agent discovers a set of relevant ‘Action‘ IDs via a function like ‘summarizeImpact‘, this batch operation allows it to retrieve the full details of each action (such as their specific ‘source_version_id‘) to build a complete and auditable report without the high performance cost of N+1 individual requests.
3.5 API as a Foundation for Agentic Reasoning
The specified actions collectively form a comprehensive yet controlled interface to the legal knowledge graph. They are intentionally designed to be low-level. An agentic framework is not expected to simply call one action to answer a user’s query. Instead, a planning module is responsible for decomposing a complex question into a sequence or a Directed Acyclic Graph (DAG) of these actions. The power of this approach is best illustrated with examples.
Example 1: Point-in-Time Retrieval Plan
A fundamental question like "What was the text of Article 6 of the Constitution in 1999?" would be decomposed into a clean, sequential plan. The agent’s goal is to deterministically find the single correct textual version for that specific date.
@@ 3.5 API as a Foundation for Agentic Reasoning
    Ground the reference: It first calls resolveItemReference(referenceText="Article 6 of the Constitution") to get the item’s canonical ID.
@@
    Find the valid version: It then calls getValidVersion, passing the item’s ID and the target timestamp ("1999-01-01T12:00:00Z"), to get the unique historical Version object.
@@
    Retrieve the text: Finally, it calls getTextForVersion, passing the retrieved version’s ID and the desired language (e.g., "pt-BR"), to get the final, correct TextUnit.
This simple chain demonstrates how a natural language query is translated into a fully deterministic and auditable retrieval plan once the initial probabilistic grounding is complete.
Example 2: Thematic Analysis
A broader, exploratory question like "Summarize the evolution of all constitutional provisions related to ’Digital Security’ since 2000" would require a thematic analysis plan that leverages the API’s powerful server-side aggregation:
  1. 1.
    First, the agent discovers the relevant conceptual node by calling searchThemes(semantic_ query="Digital Security").
  2. 2.
    Then, instead of fetching all items and their individual histories ( which would require O(N) calls where N is the number of articles in the chapter—potentially hundreds of API calls for a single constitutional chapter), the agent makes a single, powerful call to summarizeImpact, passing the theme’s id and the specified time interval.
The server performs the complex aggregation, returning a concise ImpactReport. This report contains the aggregate statistics and the identifiers of all relevant changes. If a more detailed narrative is required, the agent can then make targeted calls to getBatchActions and getBatchItems to hydrate the specific details needed for the final summary. This demonstrates how the API transforms a potentially massive client-side processing task into a single, efficient query.
Example 3: Robustness and Multilingual Fallback
The API’s atomic nature enables agents to build robust and user-friendly behaviors, such as multilingual fallback. An agent serving a user who prefers Portuguese ("pt-BR") can be programmed to handle cases where a specific text is not available in that language. Instead of simply failing, the agent can implement a graceful fallback strategy:
  1. 1.
    The agent first attempts to retrieve the text in the user’s preferred language using getTextFor Version(version_id, "pt-BR").
  2. 2.
    If this call fails (e.g., by returning a 404 error), the agent does not give up. It makes a second call to retrieve the text in a default fallback language, such as English: getTextForVersion(version_id, "en").
  3. 3.
    When presenting the result, the agent can now inform the user about the substitution: "The requested text was not available in Portuguese. Displaying the English version instead."
This pattern, enabled by the API’s clear separation of concerns, demonstrates how composable primitives can be used to build sophisticated, resilient, and transparent agentic behaviors that go beyond simple query-response loops.
3.6 Formal OpenAPI Specification
The complete, machine-readable specification of this API is provided as an OpenAPI 3.0 YAML document, available in the project repository [^1] https://github.com/hmartim/sat-graph-api. The OpenAPI specification includes: Complete request/response schemas for all actions; Example requests and responses; Detailed error response specifications; Authentication requirements; Rate limiting policies.
Developers can use this specification to generate client SDKs, test suites, and interactive documentation using standard OpenAPI tooling.
4 Use Cases and Application
To demonstrate the practical utility and expressive power of the Canonical Action API, this section presents three representative use cases. These scenarios are designed to highlight how an agentic framework can leverage the API to answer complex legal questions that are intractable for standard RAG systems. Each use case follows a pattern: a user’s query is presented, the challenge for conventional systems is outlined, and an executable plan composed of our API actions is detailed.
4.1 Use Case 1: Point-in-Time Comparison and Causal Pinpointing
@@
A standard RAG system would likely retrieve multiple, conflicting versions of Article 6 from its index, unable to deterministically isolate the versions immediately preceding and succeeding a specific conceptual change. It lacks the causal understanding to link the introduction of a word to a specific legislative event.
@@
The agent’s plan is to find the specific Action that introduced the change, which inherently contains the identifiers for the "before" and "after" states.
@@
    Identify the Item: The agent first grounds the query by resolving the textual reference to its canonical identifier.
    candidates = resolveItemReference(reference_text="Article 6 of the Constitution")
    target_item_id = candidates[0].id // Agent proceeds with the top candidate
@@
    Retrieve the Full Historical Lineage: The agent fetches the complete, time-ordered history of all legislative events that affected the item.
    history = getItemHistory(item_id=target_item_id) // Returns a list[Action]
@@
    Pinpoint the Pivotal Action (Agent Logic): The agent must now find the first Action in the history whose resulting text contains the keyword "moradia". To do this efficiently, it first collects all produces_version_ids from the ‘history‘ list. It then makes a single getBatchTexts call to retrieve all corresponding texts at once. By iterating through these texts, it can identify the pivotal Action responsible for the change. By iterating through the returned texts and their associated source_node_ids, the agent can map the first occurrence of the keyword back to the pivotal Action.
@@
    Compare the "Before" and "After" States: The pivotal Action object itself contains the direct references to the state it terminated and the state it created. The agent uses these to get a granular difference report.
    // pivotal_action is the Action object found in the previous step
    version_id_before = pivotal_action.terminates_version_id
    version_id_after = pivotal_action.produces_version_id
    diff_report = compareVersions(
    version_id_a=version_id_before,
    version_id_b=version_id_after
@@
The agent receives a structured TextDiffReport. This verifiable data, along with details from the pivotal_action itself (like its date and source_version_id), is passed to an LLM. The model can then generate a precise, auditable, and natural-language summary: "The right to ’housing’ (moradia) was introduced into Article 6 by an action on [date from action], which terminated the validity of version [ID before] and created version [ID after]. The exact change was the addition of the word ’moradia’…". This answer is precise, verifiable, and causally grounded, demonstrating a capability far beyond standard RAG systems.
4.2 Use Case 2: Causal Lineage Tracing for Legal Audit
User Query: "A legal case from 2012 references Article 227, paragraph 4 of the Constitution. Which specific law introduced this paragraph, and is its text still the same today?"

### v2

@@ 3.3.5 Introspection & Metadata Primitive
    • Discussion: This primitive enables an agent to be aware of its own knowledge boundaries. It should be used to proactively validate a user’s temporal query before attempting retrieval, allowing for graceful handling of requests for dates outside the known history and improving the overall user experience.
  2. 17.
@@
    • Discussion: This primitive provides the canonical list of values that can be used in the action_ types filter of other functions like getActionsBySource. It allows a client application to dynamically build filtering interfaces without hardcoding ontology-specific values.
  4. 19.
@@
    • Discussion: This primitive serves as the primary and most efficient entry point for an agent or user interface to explore the thematic ontology. By returning the full root objects, it allows a client to render the top-level categories without requiring subsequent API calls.
@@ 3.3.6 Architectural Rationale and Modeling Principles
Our framework is designed to handle various types of legislative state transitions with precision. This is fully represented in the Action data model, where the optionality of its properties allows an agent to unambiguously distinguish between different event types:
  • A Creation event only has a produces_version_id.
  • A pure Repeal or Revocation event might only have a terminates_version_id.
  • A standard Amendment has both.
This removes any guesswork from the agent’s causal analysis. A recommended modeling practice, for instance, is to represent a revocation not as a pure repeal, but as an amendment that creates a new version with the text "(Revoked)", thus maintaining a continuous version timeline. The API’s flexibility supports both patterns.
Discussion: Atomic Primitives and Emergent Event Cohesion
A core design principle of the SAT-Graph is the atomicity of causal events. Our model defines an Action as the most granular unit of change possible to ensure maximum auditability. A key consequence is that a single, complex legislative event—such as the enactment of "Constitutional Amendment No. 99"—is not represented by a monolithic node. Instead, its total effect is captured by an emergent collection of many individual Action nodes, whose cohesion is established by their respective source_version properties all tracing back to the same source Work.
This design provides the best of both worlds for an intelligent agent. It can use a single, high-level API primitive (getActionsBySource) to efficiently grasp the full scope of a legislative event:
@@
This powerful aggregation empowers the agent to perform high-level impact analysis, while the atomic nature of each returned Action preserves the granular verifiability needed for a trustworthy audit trail. The agent can see both the forest and the trees, deterministically.
@@
This flexibility is directly exposed through the API, particularly via the aspects parameter in functions like searchTextUnits. All TextUnits, regardless of their source entity or aspect, are candidates for embedding and semantic retrieval. This design creates a rich, multi-faceted search space. It empowers an agent with holistic perception, allowing it to discover relevant information based not just on the law’s formal content, but also on its contextual metadata and conceptual definitions in a single, unified query. This prevents the agent from developing ’tunnel vision’, ensuring that its initial, probabilistic discovery phase is as comprehensive as possible before it commits to a deterministic plan.
@@ 3.4 Practical Implementation Considerations
While the atomic primitives specified above provide the logical foundation for auditability, a production-grade API must also enable agents to execute complex reasoning plans efficiently. This section details practical considerations, most notably batch operations, which are essential for making sophisticated, multi-step agentic workflows viable in a real-world setting.
@@
An agent’s reasoning plan is not a single query, but a sequence of primitives. Scenarios like reconstructing the state of an entire constitutional chapter or analyzing a long causal history would require dozens or even hundreds of atomic calls. To make such complex plans executable, the API must provide batch-enabled versions of its core deterministic primitives. These are not mere optimizations; they are fundamental tools that allow an agent to gather large sets of evidence efficiently without flooding the network with N+1 individual requests. Key batch primitives include:
  1. 20.
@@
    • Discussion: This is a fundamental utility primitive used after any discovery process. When an agent obtains a list of ids from functions like ‘searchItems‘ or ‘getItemHierarchy‘, this batch primitive allows it to hydrate those references into full objects to access their labels, types, or hierarchical relationships without incurring the high cost of N+1 individual requests.
  2. 21.
    getBatchValidVersions(item_ids: list[ID], timestamp: date-time) $\rightarrow$ list[Version]
@@
    • Discussion: This primitive is essential for efficiently reconstructing the state of a complex hierarchical item at a specific point in time. It is designed to be used directly with the output of ‘getItemHierarchy‘, avoiding the "N+1 query problem" and drastically improving performance.
  3. 22.
    getBatchTextUnits(requests: list[{source_node_type: string,
    source_node_id: ID, language: string, aspects?: list[string]}])
    $\rightarrow$ list[TextUnit]
    • Description: A highly flexible, batch-optimized primitive that retrieves multiple TextUnits for a variety of source entities in a single request.
@@
      • requests: A list of request objects. Each object in the list must specify:
          source_node_type: The type of the source entity (e.g., "Version", "Theme").
          source_node_id: The unique identifier of the source entity.
          language: The desired language code (e.g., "pt-BR").
          aspects: (Optional) A list of textual aspects to retrieve (e.g., ["canonical", "summary"]).
    • Returns: A single list containing all the requested TextUnit objects, preserving the order of the original request list.
    • Discussion: This powerful primitive is critical for agent efficiency. After executing a complex plan and gathering a diverse set of IDs (some Versions, some Themes), an agent can use this single batch call to hydrate all necessary textual context at once, dramatically reducing network latency.
  4. 23.
@@
    • Discussion: This primitive is essential for the "hydration" step of analytical queries. For example, after an agent retrieves all primitives caused by a specific law via getActionsBySource, it might only receive a list of IDs. This batch operation allows it to retrieve the full details of each action (such as their specific terminated and produced versions) to build a complete and auditable impact report without the high performance cost of N+1 individual requests.
3.5 The API as a Grammar for Agentic Reasoning
The specified primitives are not merely a list of endpoints; they are a formal grammar for agentic reasoning. Collectively, they form a comprehensive yet controlled interface to the temporal knowledge graph. By design, they are low-level primitives. An agent is not expected to answer a user’s query with a single primitive. Instead, its planning module must act as a translator, decomposing a complex question into a Directed Acyclic Graph (DAG) of these primitives. The resulting DAG is the agent’s auditable reasoning plan. The power of this architectural pattern is demonstrated in the following section through a series of practical use cases.
3.6 Formal OpenAPI Specification
The complete, optimized and machine-readable specification of this API is provided as an OpenAPI 3.0 YAML document, available in the project repository [^2] Further examples, detailed execution plans, and the full OpenAPI specification are available in the project’s public repository: https://github.com/hmartim/sat-graph-api.. The OpenAPI specification includes: complete request/response schemas for all primitives; detailed examples of context recovery planning etc. Developers can use this specification to generate client SDKs, test suites, and interactive documentation using standard OpenAPI tooling.
4 Use Cases and Application
To demonstrate the practical utility and expressive power of the Canonical Primitive API, this section presents representative use cases. These scenarios are designed to highlight how an agent’s planning module can compose the API’s atomic primitives into verifiable reasoning plans to answer legal questions that are intractable for standard RAG systems. Each use case follows a pattern: a user’s natural language query is presented, and an executable plan composed of our API primitives is detailed.
4.1 Use Case 1: Deterministic Point-in-Time Retrieval
User Query: "What was the text of Article 6 of the Brazilian Constitution on May 20th, 2001?"
Challenge
This seemingly simple question is impossible for a standard, temporally-naive RAG system to answer reliably. A flat-text index would likely contain multiple, conflicting versions of "Article 6" from different points in time. The system lacks the temporal awareness to understand what "on May 20th, 2001" means, and cannot guarantee that the version it retrieves is the one that was legally in force on that specific date. It would likely return the most recent or most semantically similar version, leading to a factually incorrect and misleading answer. [^3] One might argue that a RAG system could be enhanced by indexing each text chunk with temporal metadata (e.g., ‘valid_from‘, ‘valid_to‘) to be used in a pre- or post-retrieval filter. While this improves precision, the core retrieval mechanism remains a probabilistic semantic search over all chunks that satisfy the temporal filter. There would be no guarantee of retrieving the single, canonically correct version, especially if multiple text chunks from different but overlapping sources were valid on the same date. Our approach, in contrast, uses a deterministic, primary-key-based lookup (getValidVersion) on a specific, pre-resolved item, eliminating retrieval uncertainty entirely.
Agent Execution Plan
The agent’s plan is a clean, sequential execution designed to isolate the single, correct version of the text. The entire process becomes deterministic after the initial grounding step.
@@ 4.1 Use Case 1: Deterministic Point-in-Time Retrieval
    Ground the Reference: The agent first translates the ambiguous natural language reference into a formal, canonical ID. The added specificity "Brazilian Constitution" helps the resolveItem Reference primitive to find the correct entity with high confidence. The agent planner also knows that the canonical text of an article resides in its caput.
    resolveItemReference(
    reference_text="Article 6, caput of the Brazilian Constitution"
@@
    Find the Valid Version: Using the now-stable ID, the agent makes a deterministic call to getValid Version, passing the item’s ID and the precise target timestamp to retrieve the unique historical Version object that was valid on that day.
    getValidVersion(
    item_id="...",
    timestamp="2001-05-20T12:00:00Z"
@@
    Retrieve the Ground-Truth Text: Finally, with the unique version ID, the agent retrieves the exact, unaltered text. The context "Brazilian" informs the choice of language. This final step is also fully deterministic.
    getTextForVersion(
    version_id="...",
    language="pt-BR"
Synthesized Outcome
The agent receives a single, verifiable TextUnit object. This structured data—the text itself, along with its version ID and validity interval—is then passed to an LLM. The model can now generate a precise and trustworthy answer: "On May 20th, 2001, the text of Article 6 of the Brazilian Constitution was: ’…’". The answer is not a probabilistic guess but a factual statement, fully grounded in a deterministically retrieved and auditable piece of evidence.
4.2 Use Case 2: Causal Pinpointing and Version Comparison
@@
This query is impossible for a standard RAG system. It requires not just retrieving text, but understanding causality—linking a conceptual change (the introduction of "housing") to a specific, discrete legislative event. A standard system lacks a causal model; it cannot deterministically find the pivotal amendment, nor can it isolate the precise "before" and "after" states that this event created. It can only search for the keyword, returning a list of contextually jumbled and unverified text chunks.
@@
The agent’s plan is to perform a historical investigation. It first retrieves the complete lineage of the article and then analyzes it to pinpoint the exact Action responsible for the change. By design, the final inspection of the texts is performed by the agent itself, making this analytical step an explicit part of its auditable reasoning plan, rather than hiding it within an opaque server-side function. This pivotal Action itself contains the direct links to the "before" and "after" versions.
@@
    Ground the Reference: As in the previous use case, the agent first grounds the query by resolving the textual reference to its canonical identifier.
    resolveItemReference(
    reference_text="Article 6, caput of the Brazilian Constitution"
@@
    Retrieve the Full Historical Lineage: The agent fetches the complete, time-ordered history of all legislative events that have ever affected the item. This provides the full evidentiary record for its investigation.
    getItemHistory(item_id="...") // Returns a list[Action]
@@
    Pinpoint the Pivotal Action: The agent must now identify the specific Action that introduced the keyword. To do this efficiently, it avoids N+1 calls by first collecting all produces_versions from the history and making a single getBatchTexts call. It then programmatically inspects the returned texts to find the first occurrence of "moradia" and identifies the corresponding pivotal Action.
@@
    Compare the "Before" and "After" States: The pivotal Action object itself contains the direct, verifiable references to the state it terminated and the state it created. The agent uses these IDs to request a granular, server-side difference report.
    // pivotal_action is the Action found in the previous step
    compareVersions(
    version_id_a=pivotal_action.terminates_version,
    version_id_b=pivotal_action.produces_version
@@
The agent receives a structured TextDiffReport. This verifiable data, along with critical metadata from the pivotal_action object itself (such as its effective date and the ID of the provision that authorized it), is passed to an LLM. The model can now generate a precise, auditable, and natural-language summary: "The right to ’housing’ (moradia) was introduced into Article 6 by Article 1 of Constitutional Amendment No. 26, effective on February 14th, 2000. This action [ID before] and [ID after] The exact textual change was the addition of the word ’moradia’ after ’lazer’." This answer is not just correct; it is causally grounded, demonstrating a level of explainability far beyond standard RAG systems.
4.3 Use Case 3: Thematic and Hierarchical Impact Analysis
User Query: "Summarize the evolution of all provisions related to the theme ’Digital Security’ since 2000."